### v4

# Jacobians with automorphisms of prime order

Abstract.
In this paper we study principally polarized abelian varieties $(X,\lambda)$ that admit an automorphism $\delta$ of prime order $p>2$. It turns out that certain natural conditions on the multiplicities of the action of $\delta$ on $\Omega^{1}(X)$ do guarantee that those polarized varieties are not canonically polarized jacobians of curves.
The author was partially supported by Simons Foundation Collaboration grant # 585711 and the Travel Support for Mathematicians Grant MPS-TSM-00007756 from the Simons Foundation. This paper was finished in August-September 2025 during his stay at Max-Planck-Institut für Mathematik (Bonn, Germany), whose hospitality and support are gratefully acknowledged.
1. Introduction
The work of Clemens and Griffiths [3] on intermediate jacobians of threefolds and their applications to the Lüroth problem increased an interest in the classical question - how to find out that a given principally polarized $g$-dimensional complex abelian variety $(X,\lambda)$ is not isomorphic to the canonically polarized jacobian $({\mathcal{J}},\Theta)$ of a smooth irreducible projective curve $\mathcal{C}$} of genus $g$. In this paper we address this question in the case when $(X,\lambda)$ admits an additional symmetry - an automorphism $\delta$ of prime period $p>2$ that satisfies the $p$th cyclotomic equation. We give our answer in terms of the multiplicity function $\mathbf{a}_{X}=\mathbf{a}_{X,\delta}$, which assigns to each $h\in({\mathbb{Z}}/p{\mathbb{Z}})^{*}$ the multiplicity $\mathbf{a}_{X}(h)$ of the eigenvalue $\zeta_{p}^{h}$ of the linear operator in the space $\Omega^{1}(X)$ of differentials of the first kind on $X$ induced by $\delta$. Namely, we describe explicitly all integer-valued (we call them strongly admissible) functions $f:({\mathbb{Z}}/p{\mathbb{Z}})^{*}\to{\mathbb{Z}}_{+}$ that enjoy the following property: there exists a triple $({\mathcal{J}},\Theta,\delta)$ as above such that $f=\mathbf{a}_{{\mathcal{J}},\delta}$.
It turns out that not every function $\mathbf{a}_{X,\delta}$ coincides with some $\mathbf{a}_{{\mathcal{J}},\delta}$. For example, if $p=3$ then there are precisely $(g+1)$ functions of type $\mathbf{a}_{X,\delta}$; however, there are approximately only $g/3$ functions of type $\mathbf{a}_{{\mathcal{J}},\delta}$ (Section 4, see also [19]).
The paper is organized as follows. In Section 2 we study canonically polarized jacobians $({\mathcal{J}},\Theta)$ of curves $\mathcal{C}$} such that there is automorphism $\delta\in\mathrm{Aut}({\mathcal{J}},\Theta)$ with properties described above. It turns out that $\delta$ is induced by an automorphism $\delta_{{\mathcal{C}}}$ of $\mathcal{C}$} of order $p$. It turns out that the number of fixed points of $\delta_{{\mathcal{C}}}$ is $\frac{2g}{p-1}+2$ and each of these points $P$ is nondegenerate, i.e., its index $\epsilon_{P}\neq 1$. This gives rise to the integer-valued function $\mathbf{b}:({\mathbb{Z}}/p{\mathbb{Z}})^{*}\to{\mathbb{Z}}_{+}$ that assigns to each $h\in({\mathbb{Z}}/p{\mathbb{Z}})^{*}$ the number of fixed points of $\delta_{{\mathcal{C}}}$ with index $\zeta_{p}^{h}$. Our main result (Theorem 2.4) expresses explicitly the function $\mathbf{a}_{{\mathcal{J}},\delta}$ in terms of the function $\mathbf{b}$, which imposes the restrictions on the function $\mathbf{a}_{{\mathcal{J}},\delta}$. In Section 3 we prove Theorem 3.1, which, in particular, asserts that the necessary condition for a function $f:({\mathbb{Z}}/p{\mathbb{Z}})^{*}\to{\mathbb{Z}}_{+}$ to coincide with some $\mathbf{a}_{{\mathcal{J}},\delta}$ imposed by Theorem 3.1 is actually sufficient. In Section 4 we discuss in detail the case $p=3$.
Acknowledgements I am grateful to George Andrews for a very informative letter about partitions.
2. Principally polarized abelian varieties with automorphisms
We write $\mathbb{Z}$$_{+}$ for the set of nonnegative integers, $\mathbb{Q}$} for the field of rational numbers and $\mathbb{C}$} for the field of complex numbers. We have
$\mathbb{Z}$$_{+}\subset{\mathbb{Z}}\subset{\mathbb{Q}}\subset$$\mathbb{R}$$\subset{\mathbb{C}}$
where $\mathbb{Z}$} is the ring of integers and $\mathbb{R}$} is the field of real numbers. If $B$ is a finite (may be, empty) set then we write #(B)\#(B) for its cardinality. Let $p$ be an odd prime and $\zeta_{p}\in{\mathbb{C}}$ a primitive (complex) $p$th root of unity. It generates the multiplicative order $p$ cyclic group $\mu_{p}$ of $p$th roots of unity. We write ${\mathbb{Z}}[\zeta_{p}]$ and ${\mathbb{Q}}(\zeta_{p})$ for the $p$th cyclotomic ring and the $p$th cyclotomic field respectively. We have
$\zeta_{p}\in\mu_{p}\subset{\mathbb{Z}}[\zeta_{p}]\subset{\mathbb{Q}}(\zeta_{p})\subset{\mathbb{C}}.$
Let $g\geq 1$ be an integer and $(X,\lambda)$ a principally polarized $g$-dimensional abelian variety over the field $\mathbb{C}$} of complex numbers, $\delta$ an automorphism of $(X,\lambda)$ that satisfies the cyclotomic equation
(1) $\sum_{j=0}^{p-1}\delta^{j}=0\in\mathrm{End}(X).$
In other words, $\delta$ is a periodic automorphism of order $p$, whose set of fixed points is finite. This gives rise to the embeddings
${\mathbb{Z}}[\zeta_{p}]\hookrightarrow\mathrm{End}(X),1\mapsto 1_{X},\ \zeta_{p}\mapsto\delta,$
${\mathbb{Q}}(\zeta_{p})\hookrightarrow\mathrm{End}^{0}(X),1\mapsto 1_{X},\ \zeta_{p}\mapsto\delta.$
Since the degree ${\mathbb{Q}}(\zeta_{p}):{\mathbb{Q}}]=p-1$, it follows from [13, Ch. 2, Prop. 2] (see also [12, Part II, p. 767]) that
(2) $(p-1)\mid 2g.$
By functoriality, ${\mathbb{Q}}(\zeta_{p})$ acts on the $g$-dimensional complex vector space $\Omega^{1}(X)$ of differentials of the first kind on $X$. This endows $\Omega^{1}(X)$ with the structure of a
${\mathbb{Q}}(\zeta_{p})\otimes_{{\mathbb{Q}}}{\mathbb{C}}$-module. Clearly,
${\mathbb{Q}}(\zeta_{p})\otimes_{{\mathbb{Q}}}{\mathbb{C}}=\oplus_{j=1}^{p-1}{\mathbb{C}}$
where the $j$th summand corresponds to the field embedding ${\mathbb{Q}}(\zeta_{p})\hookrightarrow{\mathbb{C}}$ that sends $\zeta_{p}$ to $\zeta_{p}^{j}$. So, ${\mathbb{Q}}(\zeta_{p})$ acts on $\Omega^{1}(X)$ with multiplicities $a_{j}$ ($j=1,\dots p-1$). Clearly, all $a_{j}$ are non-negative integers and
(3) $\sum_{j=1}^{p-1}a_{j}=g.$
In addition,
(4) $a_{j}+a_{p-j}=\frac{2g}{p-1}\ \forall j=1,\dots p-1;$
this is a special case of a general well known result about endomorphism fields of complex abelian varieties: see, e.g., [9, p. 84].
We may view $\{a_{j}\}$ as a nonnegative integer-valued function
$\mathbf{a}=\mathbf{a}_{X}=\mathbf{a}_{X,\delta}$
on the finite cyclic group $G=({\mathbb{Z}}/p{\mathbb{Z}})^{*}$ of order (p1)(p-1) where
(5) $\mathbf{a}_{X}(j\bmod p)=\mathbf{a}(j\bmod p)=a_{j}\ (1\leq j\leq p-1);\quad\sum_{h\in G}\mathbf{a}_{X}(h)=g.$
The group $G$ contains two distinguished elements, namely, the identity element $1\bmod p$ and the only element $(-1)\bmod p=(p-1)\bmod p$ of order $2$. If $h$ is an element of $G$ then we write h-h for the product $(-1)\bmod p\cdot h$ in $G$. If $h=j\bmod p$ then $h=(p-j)\bmod p$. In light of (4),
(6) $\mathbf{a}_{X}(h)+\mathbf{a}_{X}(-h)=\frac{2g}{p-1}\ \forall h\in G.$
For each $h=j\bmod p\in G$ we write
$\zeta_{p}^{h}:=\zeta_{p}^{j},$
which is a primitive $p$th root of unity that does not depend on the choice of $j$.
Definition 2.1.
Let $p$ be an odd prime and $g$ a positive integer such that p1p-1 divides $2g$.
Let $f:G\to{\mathbb{Z}}_{+}$ be a nonnegative integer-valued function.
  • (i)
    We say that $f$ is well rounded of degree $g$ if
    $f(h)+f(-h)=\frac{2g}{p-1}\ \forall h\in G.$
  • (ii)
    We say that $f$ is admissible of degree $g$ if there exist a principally polarized $g$-dimensional complex abelian variety $(X,\lambda)$ with an automorphism $\delta\in\mathrm{Aut}(X,\lambda)$ of period $p$ such that (1) holds and
    $f=\mathbf{a}_{X}.$
  • (iii)
    We say that $f$ is strongly admissible of degree $g$ if it is admissible of degree $g$ and one may choose the corresponding $(X,\lambda,\delta)$ in such a way that $(X,\lambda)$ is the canonically polarized jacobian of a smooth irreducible connected projective curve of genus $g$.
Remark 2.2.
  1. (i)
    In light of (6), our $\mathbf{a}=\mathbf{a}_{X}$ is well rounded. In other words, every admissible function is well rounded.
  2. (ii)
    The number of well rounded functions (for given $g$ and $p$) is obviously
    $\left(\frac{2g}{p-1}+1\right)^{(p-1)/2}.$
  3. (iii)
    Let $f:G\to{\mathbb{Z}}_{+}$ be a well rounded function of degree $g$. Then
    $\bar{f}:G\to{\mathbb{Z}}_{+},\ h\mapsto f(-h)=\frac{2g}{p-1}-f(h)$
    is also well rounded of degree $g$.
    In addition, if $f$ is admissible (resp. strongly admissible) then $\bar{f}$ is also admissible (resp. strongly admissible). Namely, let $(X,\lambda)$ be a principally polarized abelian variety, $\delta\in\mathrm{Aut}(X,\lambda)$ an automorphism of period $p$ that satisfies the $p$th cyclotomic equation (1) in $\mathrm{End}(X)$ such that $f$ coincides with the corresponding multiplicity function $\mathbf{a}_{X,\delta}$. Then $\delta^{-1}$ is also an automorphism of $\mathrm{Aut}(X,\lambda)$ of period $p$ that satisfies the $p$th cyclotomic equation and such that
    $\bar{f}=\mathbf{a}_{X,\delta^{-1}}.$
    So, $\bar{f}$ is admissible. In particular, if $(X,\lambda)=(\mathcal{J}(\mathcal{C}),\Theta)$ is the canonically polarized jacobian of a curve $\mathcal{C}$ then
    $\bar{f}=\mathbf{a}_{\mathcal{J}(\mathcal{C}),\delta^{-1}}$
    is strongly admissible.
Example 2.3.
Let $p=3$ and $E$ an elliptic curve over $\mathbb{C}$} with complex multiplication by ${\mathbb{Z}}[\zeta_{3}]$. We may take as $E$ the smooth projective model of $y^{2}=x^{3}-1$ where $\gamma_{3}$ acts on $E$ by
$\delta_{E}:(x,y)\mapsto(x,\gamma_{3}y).$
Clearly, $\delta_{E}$ satisfies the 3rd cyclotomic equation and respects the only principal polarization on $E$.
Let $g$ be a positive integer, and $f(1)$ and $f(2)$ are nonnegative integers, whose sum is $g$. Let us put
$Y_{1}=E^{f(1)},Y_{2}=E^{f(2)},\ Y=Y_{1}\times Y_{2}.$
Let $\lambda_{Y}$ be the principal polarization on $Y$ that is the product of $g$ pull-backs of the polarization on $E$. Let us consider the automorphism $\delta_{3}$ of $Y$ that acts (diagonally) as $\delta_{E}$ on $Y_{1}=E^{f(1)}$ and as $\delta_{E}^{2}=\delta_{E}^{-1}$ on $Y_{2}=E^{f(2)}$. Clearly, $\delta_{3}$ satisfies the 3rd cyclotomic equation and respects $\lambda_{Y}$. It is also clear that
$\mathbf{a}_{Y}(1)=f(1),\quad\mathbf{a}_{Y}(1)=f(2).$
In other words, if $p=3$ then every well rounded function of degree $g$ is admissible. We will see (Example 4.1 below) that not every such function is strongly admissible.
We will also need the function
(7) $\mathbf{j}:G=({\mathbb{Z}}/p{\mathbb{Z}})^{*}\to{\mathbb{Z}},\ (j\bmod p)\mapsto j\ (1\leq j\leq p-1).$
Clearly,
(8) $\mathbf{j}(h_{1}h_{2})\equiv\mathbf{j}(h_{1})\mathbf{j}(h_{2})\bmod p\ \forall h_{1},h_{2}\in G.$
Recall that if $f_{1}(h)$ and $f_{2}(h)$ are complex-valued functions on $G$ then its convolution is the function $f_{1}*f_{2}(h)$ on $G$ defined by
(9) $f_{1}*f_{2}(h)=\frac{1}{p-1}\sum_{u\in G}f_{1}(u)f_{2}(u^{-1}h).$
Theorem 2.4.
Suppose that $(X,\lambda)$ is the jacobian of a smooth projective irreducible genus $g$ curve $\mathcal{C}$} with canonical principal polarization. Then there exists a nonnegative integer-valued function
$\mathbf{b}:G=({\mathbb{Z}}/p{\mathbb{Z}})^{*}\to{\mathbb{Z}}_{+}\subset{\mathbb{C}}$
such that
(10) $\sum_{h\in G}\mathbf{b}(h)=\frac{2g}{p-1}+2,\quad\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})\in p{\mathbb{Z}},$
(11) $\mathbf{a}(v):=\mathbf{a}_{X}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}(-v)-1\ \forall v\in G.$
Example 2.5.
Let us describe explicitly the case $g=1$. Then $X$ is an elliptic curve. It follows from (2) that (p1)(p-1) divides $2$ and therefore
[…]
where
$\bar{1}=1\bmod 3,\quad\bar{2}=2\bmod 3=-\bar{1}.$
We also have
$\mathbf{j}(\bar{1})=1,\quad\mathbf{j}(\bar{2})=2.$
Then either
(12) $\mathbf{a}_{X}(\bar{1})=1,\quad\mathbf{a}_{X}(\bar{2})=0$
or
(13) $\mathbf{a}_{X}(\bar{1})=0,\quad\mathbf{a}_{X}(\bar{2})=1.$
In both cases we need to find nonnegative integers $\mathbf{b}(\bar{1})$ and $\mathbf{b}(\bar{2})$ such that
$\mathbf{b}(\bar{1})+\mathbf{b}(\bar{2})=\frac{2\cdot 1}{3-1}+2=3;$
$\frac{1}{3}\left(\mathbf{b}(\bar{1})+2\cdot\mathbf{b}(\bar{2})\right)-1=\frac{1}{3}\left(1\cdot\mathbf{b}(\bar{1})+2\cdot\mathbf{b}(\bar{2})\right)-1=\mathbf{a}_{X}(-\bar{1})=\mathbf{a}_{X}(\bar{2});$
$\frac{1}{3}\left(\mathbf{b}(\bar{2})+2\mathbf{b}(\bar{1})\right)-1=\frac{1}{3}\left(1\cdot\mathbf{b}(\bar{2})+2\cdot\mathbf{b}(\bar{1})\right)-1=\mathbf{a}_{X}(-\bar{2})=\mathbf{a}_{X}(\bar{1}).$
In the case (12), the desired $\mathbf{b}$ is given by the formulas
$\mathbf{b}(\bar{1})=0,\quad\mathbf{b}(\bar{2})=3.$
In the case (13), the desired $\mathbf{b}$ is given by the formulas
$\mathbf{b}(\bar{1})=3,\quad\mathbf{b}(\bar{2})=0.$
Proof of Theorem 2.4.
In light of Example 2.5, we may assume that $g>1$. We may assume that $(X,\lambda)=({\mathcal{J}}({\mathcal{C}}),\Theta)$ where $\mathcal{C}$} is an irreducible smooth projective genus $g$ curve, ${\mathcal{J}}({\mathcal{C}})$ its jacobian with canonical principal polarization $\Theta$, and
$\delta\in\mathrm{Aut}({\mathcal{J}}({\mathcal{C}}),\Theta)$
satisfies the $p$th cyclotomic equation
$\sum_{i=0}^{p-1}\delta^{i}=0\in\mathrm{End}({\mathcal{J}}({\mathcal{C}})).$
This means that $\delta^{p}=1\in\mathrm{End}({\mathcal{J}}({\mathcal{C}}))$ and ${\mathcal{J}}({\mathcal{C}})^{\delta}$ is finite. The latter implies that the homomorphism $1-\delta:{\mathcal{J}}({\mathcal{C}})\to{\mathcal{J}}({\mathcal{C}})$ is an isogeny, hence, surjective.
Identifying the curve $\mathcal{C}$} with its image $\mathrm{alb}_{P_{0}}({\mathcal{C}})\subset{\mathcal{J}}({\mathcal{C}})$ (i.e., $\mathrm{alb}(P_{0})$ is the zero 0 of the group law on ${\mathcal{J}}({\mathcal{C}})$, we have
$0=\mathrm{alb}_{P_{0}}(P_{0})\in C\subset{\mathcal{J}}({\mathcal{C}}).$
If $v\in X={\mathcal{J}}({\mathcal{C}})$ then we write $T_{v}$ for the translation map
$T_{v}:X\to X,\quad x\mapsto x+v.$
By Torelli Theorem (Weil’s variant, see [14, p. 35, Hauptsatz] applied to
$\delta^{(p+1)/2}\in\mathrm{Aut}({\mathcal{J}}(C),\Theta),$
$\exists\ \phi_{1/2}\in\mathrm{Aut}({\mathcal{C}})$, $\epsilon=\pm 1$ and $z\in{\mathcal{J}}({\mathcal{C}})$ such that
$\delta^{(p+1)/2}(P)=\epsilon\ \phi_{1/2}(P)+z\ \ \forall P\in C\subset{\mathcal{J}}({\mathcal{C}}).$
This implies that
$\delta(P)=\delta^{p+1}(P)=\left(\delta^{(p+1)/2}\right)^{2}(P)=\delta^{(p+1)/2}(\epsilon\ \phi_{1/2}(P)+z)=\epsilon\ \delta^{(p+1)/2}(\phi_{1/2}(P))+\delta^{(p+1)/2}(z)=$
$\epsilon\ (\epsilon\phi_{1/2}^{2}(P)+z)+\delta^{(p+1)/2}(z)=\epsilon^{2}\phi_{1/2}^{2}(P)+\left(\epsilon\ z+\epsilon\delta^{(p+1)/2}z\right)=\phi_{1/2}^{2}(P)+\left(\epsilon\ z+\epsilon\delta^{(p+1)/2}z\right).$
It follows that $\delta(P)=\phi_{1/2}^{2}(P)-w$ with $w=-(\epsilon\ z+\epsilon\ \delta^{(p+1)/2}z)$. Let us put
$\phi:=\phi_{1/2}^{2}\in\mathrm{Aut}(C).$
Then
$\phi(P)=\delta(P)+w=T_{w}\circ\delta(P)\quad\forall P\in{\mathcal{C}}\subset{\mathcal{J}}({\mathcal{C}}),$
i.e., the following diagramm is commutative. (Here the horisonal arrows are the inclusion map.)
[…]
Choose $u\in{\mathcal{J}}(C)$ such that $v-\delta(v)=(1-\delta)v=w$. Then
$T_{w}\circ\delta=T_{v}\circ\delta\circ T_{v}^{-1},$
i.e., $T_{w}\circ\delta$ and $\delta$ are conjugate in the group $\tilde{\mathrm{Aut}}({\mathcal{J}}({\mathcal{C}})$ of biregular automorphisms of the algebraic variety ${\mathcal{J}}({\mathcal{C}})$. In particular, $T_{w}\circ\delta$ is a periodic automorphism of order $p$; hence, $\phi^{p}$ is the identity automorphism of $\mathcal{C}$}. On the other hand, $\phi$ itself is not the identity map (see below).
It is well known that the map $\Psi:\Omega^{1}({\mathcal{J}}({\mathcal{C}}))\to\Omega^{1}({\mathcal{C}})$ induced by the inclusion map ${\mathcal{C}}\subset{\mathcal{J}}({\mathcal{C}})$ is an isomorphism of $g$-dimensional complex vector spaces. On the other hand, the linear map $(T_{u})_{\Omega}:\Omega^{1}({\mathcal{J}}({\mathcal{C}}))\to\Omega^{1}({\mathcal{J}}({\mathcal{C}}))$ induced by any translation $T_{u}:{\mathcal{J}}({\mathcal{C}}))\to{\mathcal{J}}({\mathcal{C}}))$ is the identity map (because all global regular 1-forms on an abelian variety are translation-invariant). Hence, the linear maps $\delta_{\Omega}:\Omega^{1}({\mathcal{J}}({\mathcal{C}}))\to\Omega^{1}({\mathcal{J}}({\mathcal{C}}))$ and $(T_{w}\circ\delta)_{\Omega}:\Omega^{1}({\mathcal{J}}({\mathcal{C}}))\to\Omega^{1}({\mathcal{C}})$ induced by $\delta$ and $T_{w}\circ\delta$ respectively do coincide. This implies that the following diagramm is commutative.
[…]
It follows that
$\phi_{\Omega}=\Psi\circ\delta_{\Omega}\circ\Psi^{-1}.$
In particular, the linear operators $\phi_{\Omega}$ and $\delta_{\Omega}$ have the same trace, and $\phi_{\Omega}$ is an automorphism of order $p$. This implies that $\phi$ is not the identity map, hence, has order $p$.
Let $P\in{\mathcal{C}}$ be a fixed point of $\phi$. Then $\phi$ induces the automorphism of the corresponding (one-dimensional) tangent space $\mathcal{T}_{P}({\mathcal{C}})$, which is multiplication by a complex number $\epsilon_{P}$ that is called the index of $P$.
Lemma 2.6.
Every fixed point $P$ of $\phi$ is nondegenerate, i.e., $\epsilon_{P}\neq 1$.
Proof of Lemma 2.6.
The result is well-known. but I failed to find a proper reference (however, see [19, Lemma 1.2]) where the case $p=3$ was proven.)
Suppose that $c_{P}=1$. Let ${\mathcal{O}}_{P}$ be the local ring at $P$ and ${\mathfrak{m}}_{P}$ its maximal ideal. We write $\phi_{*}$ for the automorphism of ${\mathcal{O}}_{P}$ induced by $\phi$. Clearly, $\phi_{*}^{p}$ is the identity map. Since $\phi$ is not the identity map, there are no $\phi_{*}$-invariant local parameters at $P$. Clearly, $\phi_{*}({\mathfrak{m}}_{P})={\mathfrak{m}}_{P},\phi_{*}({\mathfrak{m}}_{P}^{2})={\mathfrak{m}}_{P}^{2}$. Since $T_{P}(C)$ is the dual of ${\mathfrak{m}}_{P}/{\mathfrak{m}}_{P}^{2}$ and $c_{p}=1$, we conclude that $\phi_{*}$ induces the identity map on ${\mathfrak{m}}_{P}/{\mathfrak{m}}_{P}^{2}$. This implies that if $t\in{\mathfrak{m}}_{P}$ is a local parameter at $t$ (i.e., its image $\bar{t}$ in ${\mathfrak{m}}_{P}/{\mathfrak{m}}_{P}^{2}$ is not zero) then
$t^{\prime}:=\sum_{k=0}^{p-1}\phi_{*}^{k}(t)\in{\mathfrak{m}}_{P}\subset{\mathcal{O}}_{P}$
is $\phi_{*}$-invariant and its image in ${\mathfrak{m}}_{P}/{\mathfrak{m}}_{P}^{2}$ equals $p\bar{t}\neq 0$. This implies that $t^{\prime}$ is a $\phi_{*}$-invariant local parameter at $P$. Contradiction.
Corollary 2.7.
The quotient $D:={\mathcal{C}}/\mu_{p}$ is a smooth projective irreducible curve. The map ${\mathcal{C}}\to D$ has degree $p$, its ramification points are exactly the images of fixed points of $\phi$ and all the ramification indices are $p$.
Lemma 2.8.
$D$ is biregularly isomorphic to the projective line.
Proof of Lemma 2.8.
The map ${\mathcal{C}}\to D$ induces, by Albanese functorialy, the surjective homomorphism of the corresponding jacobians $J({\mathcal{C}})\to J(D)$ that kills the divisors classes of the form $(Q)-(\phi(Q))$ for all $Q\in{\mathcal{C}}({\mathbb{C}})$. This implies that it kills $(1-\delta)J({\mathcal{C}})$. On the other hand, $1-\delta:J({\mathcal{C}})\to J({\mathcal{C}})$ is an isogeny. This implies that the image of $J({\mathcal{C}})$ in $J(D)$ is zero and the surjectiveness implies that $J(D)=0$. This means that the genus of $D$ is 0. ∎
Corollary 2.9.
The number $F(\phi)$ of fixed points of $\phi$ is $\frac{2g}{p-1}+2$.
Proof of Corollary 2.9.
Applying the Riemann-Hurwitz’s formula to ${\mathcal{C}}\to D$, we get
$2g-2=p\cdot(-2)+(p-1)\cdot F(\phi).$
Lemma 2.10.
Let $\phi^{*}:\Omega^{1}({\mathcal{C}})\to\Omega^{1}({\mathcal{C}})$ be the automorphism of the $g$-dimensional complex vector space $\Omega^{1}({\mathcal{C}})$ induced by $\phi$ and $\tau$ the trace of $\phi^{*}$. Then
$\tau=\sum_{j=1}^{p-1}a_{j}\zeta_{p}^{j}=\sum_{h\in G}\mathbf{a}(h)\zeta_{p}^{h}.$
Proof of Lemma 2.10.
Pick a $\phi$-invariant point $P_{0}\in{\mathcal{C}}({\mathbb{C}})$ and consider the regular map
$\mathfrak{a}:{\mathcal{C}}\to J(C),\quad Q\mapsto\mathrm{cl}((Q)-(P_{0})).$
It is well-known that $\mathfrak{a}$ induces an isomorphism of complex vector spaces
$\mathfrak{a}^{*}:\Omega^{1}(J({\mathcal{C}}))\cong\Omega^{1}({\mathcal{C}}).$
Clearly,
$\phi^{*}=\mathfrak{a}^{*}\delta^{*}{\mathfrak{a}^{*}}^{-1}$
where $\delta^{*}:\Omega^{1}(J({\mathcal{C}}))=\Omega^{1}(J({\mathcal{C}}))$ is the $\mathbb{C}$}-linear automorphism induced by $\delta$. This implies that the traces of $\phi^{*}$ and $\delta^{*}$ do coincide. Now the very definition of $a_{j}$’s implies that the trace of $\phi^{*}$ equals $\sum_{j=1}^{p-1}a_{j}\zeta_{p}^{j}$. ∎
Lemma 2.11.
Let $\zeta\in{\mathbb{C}}$ be a primitive $p$th root of unity. Then
(14) $\frac{1}{1-\zeta}=-\frac{\sum_{j=1}^{p-1}j\zeta^{j}}{p}=-\frac{\sum_{h\in G}\mathbf{j}(h)\zeta^{h}}{p}.$
Proof of Lemma 2.11.
We have
$(1-\zeta)\left(\sum_{j=1}^{p-1}j\zeta^{j}\right)=\sum_{j=1}^{p-1}\left(j\zeta^{j}-j\zeta^{j+1}\right)=\left(\sum_{j=1}^{p-1}\zeta^{j}\right)-(p-1)\zeta^{p}=(-1)-(p-1)=-p.$
End of proof of Theorem 2.4. Let $B$ be the set of fixed points of $\phi$. We know that $(B)=\frac{2g}{p-1}+2$. By the holomorphic Lefschetz fixed point formula [2, Th. 2], [4, Ch. 3, Sect. 4] (see also [8, Sect. 12.2 and 12.5]) applied to $\phi$,
(15) $1-\bar{\tau}=\sum_{P\in B}\frac{1}{1-\epsilon_{P}}$
where $\bar{\tau}$ is the complex-conjugate of $\tau$. Recall that every $\epsilon_{P}$ is a (primitive) $p$th root of unity. Now Theorem 2.4 follows readily from the following assertion.
Proposition 2.12.
Let us define for each $h\in G$ the nonnegative integer $\mathbf{b}(h)$ as the number of fixed points $P\in B\subset{\mathcal{C}}({\mathbb{C}})$ such that $\epsilon_{P}=\zeta_{p}^{h}$. Then
(16) $\sum_{h\in G}\mathbf{b}(h)=F(\phi)=\frac{2g}{p-1}+2.$
and
(17) $\mathbf{a}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}(-v)-1\ \forall v\in G.$
In particular,
${\mathbb{Z}}\ni 1+\mathbf{a}(-1\bmod p)=\frac{1}{p}\left(\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})\right).$
Proof of Proposition 2.12.
The equality (16) is obvious. Let us prove (17). Combining (15) with Lemma 2.11 (applied to $\zeta=\zeta_{p}^{h}$) and Lemma 2.10, we get
$1-\sum_{h\in G}\mathbf{a}(h)\zeta_{p}^{-h}=\sum_{u\in G}\mathbf{b}(u)\frac{1}{1-\zeta_{p}^{u}}=\frac{-1}{p}\left(\sum_{u\in G}\mathbf{b}(u)\left(\sum_{h\in G}\mathbf{j}(h)\zeta_{p}^{hu}\right)\right)=$
$\frac{-1}{p}\sum_{v\in G}\left(\sum_{u\in G}\mathbf{b}(u)\mathbf{j}(u^{-1}v)\right)\zeta_{p}^{v}=\frac{-1}{p}\sum_{v\in G}(p-1)\mathbf{b}*\mathbf{j}(v)\zeta_{p}^{v}=\frac{-(p-1)}{p}\sum_{v\in G}\mathbf{b}*\mathbf{j}(v)\zeta_{p}^{v}$
(here we use a substitution $v=hu$). Taking into account that
$0=1+\sum_{j=1}^{p-1}\zeta_{p}^{j}=1+\sum_{v\in G}\zeta_{p}^{v},$
we obtain
$\left(\sum_{v\in G}\zeta_{p}^{v}\right)-\sum_{h\in G}\mathbf{a}(h)\zeta_{p}^{-h}=-\frac{(p-1)}{p}\sum_{v\in G}\mathbf{b}*\mathbf{j}(v)\zeta_{p}^{v}.$
Taking into account that the (p1)(p-1)-element set
$\{\zeta_{p}^{j}\mid 1\leq j\leq p-1\}=\{\zeta_{p}^{v}\mid v\in G\}$
is a basis of the $\mathbb{Q}$}-vector space ${\mathbb{Q}}(\zeta_{p})$, we get $1+\mathbf{a}(-v)=(p-1)\mathbf{b}*\mathbf{j}(v)/p$, i.e.,
$\mathbf{a}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}(-v)-1\ \forall v\in G.$
Remark 2.13.
Let us consider the function
(18) $\mathbf{j}_{0}=\mathbf{j}-\frac{p}{2}:G=({\mathbb{Z}}/p{\mathbb{Z}})^{*}\to{\mathbb{Q}},\ (j\bmod p)\mapsto j-\frac{p}{2}\ \ \text{ where }\ j=1,\dots,p-1.$
Then
(19) $\mathbf{j}_{0}(-u)=-\mathbf{j}_{0}(u)\ \forall u\in G.$
We have
$\mathbf{b}*\mathbf{j}(v)=\mathbf{b}*\mathbf{j}_{0}(v)+\frac{p}{2(p-1)}\sum_{h\in G}\mathbf{b}(h)=\mathbf{b}*\mathbf{j}_{0}(v)+\frac{p}{2(p-1)}\left(\frac{2g}{p-1}+2\right).$
This implies that
$\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}_{0}(v)+\frac{g}{p-1}+1$
and therefore
(20) $\mathbf{a}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}_{0}(-v)+\frac{g}{p-1}\ \forall v\in G.$
On the other hand, It follows from (19) that the convolution $\mathbf{b}*\mathbf{j}_{0}$ also satisfies
$\mathbf{b}*\mathbf{j}_{0}(-v)=\mathbf{b}*\mathbf{j}_{0}(v)\ \ \forall v\in G.$
This implies that
$\mathbf{a}(v)+\mathbf{a}(-v)=\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}_{0}(-v)+\frac{g}{p-1}+\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}_{0}(v)+\frac{g}{p-1}=\frac{2g}{p-1}\ \forall v\in G.$
This implies that
(21) $\mathbf{a}(v)+\mathbf{a}(-v)=\frac{2g}{p-1}.$
(Actually, we already know it, see (4).) It follows from (21) that
(22) $\mathbf{a}(v)=\frac{2g}{p-1}-\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}(v)+1\ \forall v\in G.$
Corollary 2.14.
We keep the notation and assumptions of Theorem 2.4. Let
$\mathbf{b}^{\prime}:G\to{\mathbb{C}}$
be a complex-valued function on $G$.
  • (a)
    The following two conditions are equivalent.
    1. (a1)
      $\mathbf{a}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}^{\prime}*\mathbf{j}(-v)-1\quad\forall v\in G$.
    2. (a2)
      The odd parts of functions $\mathbf{b}$ and $\mathbf{b}^{\prime}$ do coincide, i.e.,
      (23) $\mathbf{b}^{\prime}(v)-\mathbf{b}^{\prime}(-v)=\mathbf{b}(v)-\mathbf{b}(-v).$
      In addition,
      (24) $\sum_{h\in G}\mathbf{b}^{\prime}(h)=\sum_{h\in G}\mathbf{b}(h)=\frac{2g}{p-1}+2,$
  • (b)
    If $p=3$ and
    $\mathbf{a}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}^{\prime}*\mathbf{j}(-v)-1\quad\forall v\in G.$
    then
    $\mathbf{b}^{\prime}(v)=\mathbf{b}(v)\ \forall v\in G.$
Proof.
If $f:G\to{\mathbb{C}}$ is a complex-valued function on $G$ and $\chi:G\to{\mathbb{C}}^{*}$ is a character (group homomorphism) then we write
$c_{\chi}(f)=\frac{1}{p-1}\sum_{h\in G}f(h)\bar{\chi}(h)$
for the corresponding Fourier coefficient of $f$. For example, if $\chi_{0}\equiv 1$ is the trivial character of $G$ then
$c_{\chi_{0}}(f)=\frac{1}{p-1}\sum_{h\in G}f(h).$
In particular,
$c_{\chi_{0}}(\mathbf{j})=\frac{1}{p-1}\sum_{j=1}^{p-1}j=\frac{p}{2}\neq 0.$
We have
(25) $f(v)=\sum_{\chi\in\hat{G}}c_{\chi}(f)\chi(v)\ \text{ where }\hat{G}=\mathrm{Hom}(G,{\mathbb{C}}^{*}).$
Let us consider the function
$d:G\to{\mathbb{C}},\ d(v)=\mathbf{b}^{\prime}(v)-\mathbf{b}(v).$
Suppose that (a1) holds. We need to check that (a2) holds, i.e.,
$\sum_{h\in G}d(h)=0,\quad d(v)=d(-v)\ \forall v\in G,$
which means that
$c_{\chi_{0}}(d)=0$
and for all odd characters $\chi$ (i.e., characters $\chi$ of $G$ with
$\chi(-1\bmod p)=-1)$
the corresponding Fourier coefficient
$c_{\chi}(d)=0.$
It follows from (11) that $d*\mathbf{j}(-v)=0$ for all $v\in G$, i.e.,
$d*\mathbf{j}(v)=0\ \forall v\in G.$
This implies that
$0=c_{\chi}(d*\mathbf{j})=c_{\chi}(d)\cdot c_{\chi}(\mathbf{j})\quad\forall\chi\in\hat{G}$.
However, we know that $c_{\chi_{0}}(\mathbf{j})\neq 0$. On the other hand, $c_{\chi}(\mathbf{j})\neq 0$ for all odd $\chi$: it follows from [6, Chap. 16, Theorem 2] combined with [10, Ch. 9, p. 288, Th. 9.9], see also [16, p. 477]. This implies that $c_{\chi_{0}}(d)=0$ and $c_{\chi}(d)=0$ for all odd $\chi$. This proves that (a1) implies (a2).
Assume now that (a2) holds. This means that $d(v)$ is an even function, i.e.,
$d(-v)=d(v)\quad\forall v\in G$
and
$\sum_{v\in G}d(v)=-0.$
We need to prove that (a1) holds. Let us prove first that
(26) $\mathbf{a}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}^{\prime}*\mathbf{j}_{0}(-v)+\frac{g}{p-1}\ \forall v\in G.$
In light of (20), in order to prove (26), it suffices to check that
$d*\mathbf{j}_{0}(-v)=0\quad\forall v\in G,$
i.e.,
(27) $D_{v}:=\sum_{h\in G}d(h)\cdot\mathbf{j}_{0}(h^{-1}v)=0\quad\forall v\in G.$
In order to prove (27), recall that $\mathbf{j}_{0}$ is odd and $d$ is even. This implies that
$D_{v}=\sum_{h\in G}d(h)\cdot\mathbf{j}_{0}(h^{-1}v)=\sum_{h\in G}d(-h)\cdot\mathbf{j}_{0}((-h)^{-1}v)=$
$\sum_{h\in G}d(h)\cdot\mathbf{j}_{0}(-h^{-1}v)=\sum_{h\in G}d(h)\cdot\left(-\mathbf{j}_{0}(h^{-1}v)\right)=-\sum_{h\in G}d(h)\cdot\mathbf{j}_{0}(h^{-1}v)=-D_{v}.$
It follows that
$\sum_{h\in G}d(h)\cdot\mathbf{j}_{0}(h^{-1}v)=D_{v}=0,$
which proves (26).
Now taking into account that $\mathbf{j}_{0}=\mathbf{j}-p/2$, we get from (26) that
$\mathbf{a}(v)=\frac{1}{p}\sum_{h\in G}\mathbf{b}^{\prime}(h)\cdot\Big{(}\mathbf{j}(h^{-1}(-v))-p/2\Big{)}=$
$\left(\frac{1}{p}\sum_{h\in G}\mathbf{b}^{\prime}(h)\cdot\mathbf{j}(h^{-1}(-v)\right)-\frac{1}{2}\left(\sum_{h\in G}\mathbf{b}^{\prime}(h)\right)+\frac{g}{p-1}=$
$\frac{(p-1)}{p}\cdot\mathbf{b}^{\prime}*\mathbf{j}(-v)-\frac{1}{2}\left(\frac{2g}{p-1}+2\right)+\frac{g}{p-1}=\mathbf{b}^{\prime}*\mathbf{j}(-v)-1.$
So, $\mathbf{a}(v)=\mathbf{b}^{\prime}*\mathbf{j}(-v)-1$, i.e., (a1) holds. This ends the proof of (a).
Now let $p=3$. Then 2+2g/(p1)=g+22+2g/(p-1)=g+2 and $G=\{1\bmod 3,-1\bmod 3\}$. We already know that
$\mathbf{b}^{\prime}(1\bmod 3)-\mathbf{b}^{\prime}(-1\bmod 3)=\mathbf{b}(1\bmod 3)-\mathbf{b}(-1\bmod 3),$
$\mathbf{b}^{\prime}(1\bmod 3)+\mathbf{b}^{\prime}(-1\bmod 3)=g+2=\mathbf{b}(1\bmod 3)+\mathbf{b}(-1\bmod 3).$
This implies that
$\mathbf{b}^{\prime}(1\bmod 3)=\mathbf{b}(1\bmod 3),\quad\mathbf{b}^{\prime}(-1\bmod 3)=\mathbf{b}(-1\bmod 3),$
which proves (b). ∎
Remark 2.15.
If $v\in G$ then there is an integer $k_{v}$ that does not divide $p$ and such that $\mathbf{j}(vh)-k_{v}\mathbf{j}(h)$ is divisible by $p$ for all $h\in G$. Indeed, the function
$\gamma:G\to({\mathbb{Z}}/p{\mathbb{Z}})^{*},\ h=j\bmod p\mapsto\mathbf{j}(h)\bmod\ p=j\bmod p$
is a group homomorphism. Hence,
$\gamma(vh)=\gamma(v)\gamma(h)\ \forall v,h\in G.$
Let us choose an integer $k_{v}\in{\mathbb{Z}}$ such that
$k_{v}\bmod p=\gamma(v)=\mathbf{j}(h)\bmod p.$
Clearly, $p$ does not divide $k_{v}$ and
$\mathbf{j}(vh)\bmod p=\gamma(vh)=\gamma(v)\cdot\gamma(h)=(k_{v}\bmod p)\cdot\gamma(h)=(k_{v}\bmod p)\cdot(\mathbf{j}(h)\bmod p).$
This implies that $\mathbf{j}(vh)-k_{v}\mathbf{j}(h)$ is divisible by $p$ for all $v\in G$.
Corollary 2.16.
Let $c:G\to{\mathbb{Z}}$ be an integer-valued function. Then the following conditions are equivalent.
  • (i)
    $(p-1)\cdot c*\mathbf{j}(1\bmod p)=\sum_{h\in G}c(h)\mathbf{j}(h^{-1})\in p{\mathbb{Z}}$.
  • (ii)
    $(p-1)\cdot c*\mathbf{j}(v)=\sum_{h\in G}c(h)\mathbf{j}(h^{-1}v)\in p{\mathbb{Z}}\ \forall v\in G.$
Proof.
Clearly, (ii) implies (i). Suppose that (i) holds, i.e.,
$\sum_{h\in G}c(h)\mathbf{j}(h^{-1})\in p{\mathbb{Z}}.$
In order to prove that (ii) holds, we need to check that
$\sum_{h\in G}c(h)\mathbf{j}(h^{-1}v)\in p{\mathbb{Z}}\ \forall v\in G.$
Notice that in light of Remark 2.15 (applied to $h^{-1}$), if $v\in G$ then there exists $k_{v}\in{\mathbb{Z}}$ such that $\mathbf{j}(vh^{-1})-k_{v}\mathbf{j}(h^{-1})$ is divisible by $p$ for all $h\in G$. In other words, $\mathbf{j}(vh^{-1})\equiv k_{v}\mathbf{j}(h)\bmod p$. This implies that $\forall v\in G$
$\sum_{h\in G}c(h)\mathbf{j}(h^{-1}v)=\sum_{h\in G}c(h)\mathbf{j}(vh^{-1})\equiv k_{v}\sum_{h\in G}c(h)\mathbf{j}(h^{-1})\bmod p\equiv 0\bmod p.$
This ends the proof. ∎
The next assertion shows that not every well rounded function is strongly admissible.
Theorem 2.17.
Suppose that $(X,\lambda)$ is the jacobian of a smooth projective irreducible genus $g$ curve $\mathcal{C}$} with canonical principal polarization. Then
(28) $\frac{1}{p}\cdot\frac{2g}{(p-1)}-\frac{p-2}{p}\leq\mathbf{a}(v)\leq\frac{2g}{(p-1)}-\left(\frac{1}{p}\cdot\frac{2g}{(p-1)}-\frac{p-2}{p}\right)\quad\forall v\in G.$
In particular, if (p2)<2g/(p1)(p-2)<2g/(p-1) then
$1\leq\mathbf{a}(v)\leq\frac{2g}{p-1}-1\quad\forall v\in G.$
Proof.
By Theorem 2.4, there exists a nonnegative integer-valued function $\mathbf{b}:G\to{\mathbb{Z}}_{+}$ such that
$\mathbf{a}(h)=\frac{\sum_{u\in G}\mathbf{b}(u)\mathbf{j}(-u^{-1}h)}{p}-1\ \forall h\in G.$
Recall that
$\mathbf{b}(u)\geq 0,\ \ \sum_{u\in G}\mathbf{b}(u)=\frac{2g}{p-1}+2,\ \ 1\leq\mathbf{j}(v)\leq p-1.$
This implies that
$\frac{\sum_{u\in G}\mathbf{b}(u)}{p}-1\leq\mathbf{a}(h)\leq(p-1)\cdot\frac{\sum_{u\in G}\mathbf{b}(u)}{p}-1.$
This means that
$\frac{\frac{2g}{p-1}+2}{p}-1\leq\mathbf{a}(h)\leq(p-1)\cdot\frac{\frac{2g}{p-1}+2}{p}-1.$
Hence,
$\frac{1}{p}\cdot\frac{2g}{(p-1)}-\frac{p-2}{p}\leq\mathbf{a}(h)\leq\frac{2g}{(p-1)}-\left(\frac{1}{p}\cdot\frac{2g}{(p-1)}-\frac{p-2}{p}\right)\quad\forall v\in G,$
which ends the proof. ∎
3. A construction of jacobians
The following theorem may be viewed as an inverse of Theorem 2.4.
Theorem 3.1.
Let $g$ be a positive integer, $p$ an odd prime, $\zeta_{p}\in{\mathbb{C}}$ a primitive $p$th root of unity, and $G=({\mathbb{Z}}/p{\mathbb{Z}})^{*}$. Suppose that (p1)(p-1) divides $2g$. Let $\mathbf{b}:G\to{\mathbb{Z}}_{+}$ be a non-negative integer-valued function such that
  • (i)
    (29) $\sum_{h\in G}\mathbf{b}(h)=\frac{2g}{p-1}+2.$
  • (ii)
    (30) $(p-1)\cdot\mathbf{b}*\mathbf{j}(1\bmod p)=\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})\in p{\mathbb{Z}}.$
Let $\{f_{h}(x)\mid h\in G\}$ be a (p1)(p-1)-element set of mutually prime nonzero polynomials $f_{h}(x)\in{\mathbb{C}}[x]$ that enjoy the following properties.
  • (1)
    $\deg(f_{h})=\mathbf{b}(h)$ for all $h\in G$. In particular, $f_{h}(x)$ is a (nonzero) constant polynomial if and only if $\mathbf{b}(h)=0$.
  • (2)
    Each $f_{h}(x)$ has no repeated roots.
Let us consider a polynomial
$f(x)=f_{\mathbf{b}}(x)=\prod_{h\in G}f_{h}(x)^{\mathbf{j}(h^{-1})}\in{\mathbb{C}}[x]$
of degree $\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})$. Let $\mathcal{C}$} be the smooth projective model of the irreducible plane affine curve
(31) $y^{p}=f_{\mathbf{b}}(x)$
endowed with an automorphism $\delta_{{\mathcal{C}}}:{\mathcal{C}}\to{\mathcal{C}}$ induced by
$(x,y)\mapsto(x,\zeta_{p}y).$
Let $(\mathcal{J},\lambda)$ be the canonically principally polarized jacobian of $\mathcal{C}$} endowed by the automorphism $\delta$ induced by $\delta_{{\mathcal{C}}}$. Then $\mathcal{J}$ and $\delta$ enjoy the following properties.
  • (a)
    $\mathrm{dim}(\mathcal{J})=g$
  • (b)
    $\sum_{j=0}^{p-1}\delta^{j}=0$ in $\mathrm{End}(\mathcal{J})$.
  • (c)
    Let $\mathbf{a}=\mathbf{a}_{\mathcal{J}}:G\to{\mathbb{Z}}_{+}$ be the corresponding multiplicity function attached to the action of $\delta$ on the differentials of first kind on $\mathcal{J}$ defined in (5).
    Then
    (32) $\mathbf{a}_{\mathcal{J}}(v)=\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}(-v)-1=\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}_{0}(-v)+\frac{g}{p-1}=\frac{2g}{p-1}-\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}(v)+1$
    $\forall v\in G$.
Remark 3.2.
It follows from (29) and (30) combined with Corollary 2.16 that RHS of (32) is a nonnegative integer.
Proof of Theorem 3.1.
If $\alpha$ is a root of $f(x)$ then there is exactly one $h\in G$ that $\alpha$ is a root of $f_{h}(x)$; in addition, the multiplicity of $\alpha$ (viewed as a root of $f(x)$) is $\mathbf{j}(h^{-1})$, which is not divisible by $p$. This implies that $f(x)$ is not a $p$th power in the polynomial ring ${\mathbb{C}}[x]$ and even in the field of rational function ${\mathbb{C}}(x)$. It follows from theorem 9.1 of [7, Ch. VI, Sect. 9] that the polynomial
$y^{p}-f(x)\in{\mathbb{C}}(x)[y]$
is irreducible over ${\mathbb{C}}(x)$. This implies that the polynomial in two variable
$y^{p}-f(x)\in{\mathbb{C}}[x,y]$
is irreducible, because every its divisor that is a polynomial in $x$ is a constant, i.e., the affine plane curve (31) is irreducible and its field of rational functions $K$ is the field of fractions of the domain
$A={\mathbb{C}}[x,y]/(y^{p}-f(x)){\mathbb{C}}[x,y].$
Let $\mathcal{C}$} be the smooth projective model of the curve (31). Then $K$ is the field ${\mathbb{C}}({\mathcal{C}})$ of rational functions on $\mathcal{C}$}; in particular, ${\mathbb{C}}({\mathcal{C}})$ is generated over $\mathbb{C}$} by rational functions $x,y$. Let $\pi:{\mathcal{C}}\to\mathbb{P}^{1}$ be the regular map defined by rational function $x$. Clearly, it has degree $p$. Since
$\deg(\pi)=\deg(f)=\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})$
is divisible by $p$, the map $\pi$ is unramified at $\infty$ (see [11, Sect. 4]) and therefore the set of branch points of $\pi$ coincides with the set of roots of $f(x)$, which, in turn, is the disjoint union of the sets $R_{h}$ of roots of $f_{h}(x)$. In particular, the number of branch points of $\pi$ is
$\sum_{h\in G}\deg(f_{h})=\sum_{h\in G}\mathbf{b}(h)=\frac{2g}{p-1}+2.$
Clearly, $\pi$ is a Galois cover of degree $p$, i.e., the field extension
${\mathbb{C}}({\mathcal{C}})/{\mathbb{C}}(\mathbb{P}^{1})={\mathbb{C}}({\mathcal{C}})/{\mathbb{C}}(x)$
is a cyclic field extension of degree $p$. In addition, the cyclic Galois group $\mathrm{Gal}({\mathbb{C}}({\mathcal{C}})/{\mathbb{C}}(\mathbb{P}^{1}))$ is generated by the automorphism $\delta_{{\mathcal{C}}}:{\mathcal{C}}\to{\mathcal{C}}$ induced by
$\delta_{{\mathcal{C}}}:{\mathcal{C}}\to{\mathcal{C}},\ (x,y)\mapsto(x,\zeta_{p}y).$
It follows from the Riemann-Hurwitz formula (see [11, Sect. 4]) that the genus of $\mathcal{C}$} is
$\frac{\left(\left(\frac{2g}{p-1}+2\right)-2\right)\left(p-1\right)}{2}=g.$
In addition, the automorphism $\delta$ of the canonically polarized jacobian $(\mathcal{J},\lambda)$ induced by $\delta_{{\mathcal{C}}}$ satisfies the $p$th cyclotomic equation
$\sum_{j=0}^{p-1}\delta^{j}=0\in\mathrm{End}(\mathcal{J})$
(see [11, p. 149]). Let $B\subset{\mathcal{C}}({\mathbb{C}})$ be the set of ramification points of $\pi$. Clearly, $B$ coincides with the set of fixed points of $\delta_{{\mathcal{C}}}$. The map $x:{\mathcal{C}}({\mathbb{C}})\to\mathbb{P}^{1}({\mathbb{C}})$ establishes a bijection between $B$ and the disjoint union of all $R_{h}$’s. Let us put
$B_{h}=\{P\in B\mid x(P)\in R_{h}\}.$
Then $B$ partitions onto a disjoint union of all $B_{h}$’s and
$(B_{h})=\deg(f_{h})=\mathbf{b}(h)\ \forall h\in G.$
Let $P\in B$. The action of $\delta$ on the tangent space to $C$ at $P$ is multiplication by a certain $p$th root of unity $\epsilon_{P}$.
Lemma 3.3.
$\epsilon_{P}=\zeta_{p}^{\mathbf{j}(h)}$ if and only if $P\in B_{h}$.
Proof of Lemma 3.3.
Clearly, if $h_{1}$ and $h_{2}$ are distinct elements of $G=({\mathbb{Z}}/p{\mathbb{Z}})^{*}$ then
$1\leq\mathbf{j}(h_{1}),\ \mathbf{j}(h_{2})\leq p-1;\quad\mathbf{j}(h_{1})\neq\mathbf{j}(h_{2}).$
Therefore $\zeta_{p}^{\mathbf{j}(h_{1})}\neq\zeta_{p}^{\mathbf{j}(h_{2})}$. Hence, in order to prove our Lemma, it suffices to check that
(33) $\epsilon_{P}=\zeta_{p}^{\mathbf{j}(h)}\ \ \text{if }\ P\in B_{h}.$
So, let $P\in B_{h}$. Then we have
$x(P)=\alpha\in R_{h},\ y(P)=0.$
Let
$\mathrm{ord}_{P}:{\mathbb{C}}(C)\twoheadrightarrow{\mathbb{Z}}$
be the discrete valuation map attached to $P$. Then one may easily check that
$\mathrm{ord}_{P}(x-\alpha)=p,\ \mathrm{ord}_{P}(x-\beta)=0\ \forall\beta\in{\mathbb{C}}\setminus\{\alpha\}.$
This implies that
$p\cdot\mathbf{j}(h^{-1})\cdot\mathrm{ord}_{P}(x-\alpha)=\mathrm{ord}_{P}(y^{p})=p\cdot\mathrm{ord}_{P}(y)$
and therefore
(34) $\mathrm{ord}_{P}(y)=\mathbf{j}(h^{-1}).$
In light of (8), there is an integer $m$ such that
$\mathbf{j}(h^{-1})\cdot\mathbf{j}(h)=1+pm.$
Combining this with (34), we obtain that
$\mathrm{ord}_{P}\left(\frac{y^{\mathbf{j}(h)}}{(x-\alpha)^{m}}\right)=\mathbf{j}(h^{-1})\cdot\mathbf{j}(h)-pm=1$
and therefore $t:=y^{\mathbf{j}(h)}/(x-\alpha)^{m}$ is a local parameter of $\mathcal{C}$} at $P$. Clearly, the action of $\delta$ multiplies $t$ by $\zeta_{p}^{\mathbf{j}(h)}$ and therefore $\epsilon_{P}=\zeta_{p}^{\mathbf{j}(h)}$, which proves the Lemma. ∎
End of Proof of Theorem 3.1. Now the desired result follows from Proposition 2.12 applied to $X=\mathcal{J},\phi=\delta$ combined with (26) and (22). ∎
4. The $p=3$ case
In this section we discuss the case when $p=3$ (see also [19, 20]).
Example 4.1.
Let $p=3$. We have seen (Example 2.3) that every well rounded function of degree $g$ is admissible and the number of such functions is $(g+1)$.
Let us list all the possibilities for strongly admissible $\mathbf{a}$ when $g$ is given. Let us identify
$G=({\mathbb{Z}}/3{\mathbb{Z}})^{*}=\{1\bmod 3,2\bmod 3\}$
with the set $\{1,2\}$ in the obvious way. We have the following conditions on $\mathbf{b}$.
$\mathbf{b}(1),\mathbf{b}(2)\in{\mathbb{Z}}_{+},\ \mathbf{b}(1)+\mathbf{b}(2)=g+2,\ \ 3\ \text{divides}\ \mathbf{b}(1)+2\mathbf{b}(2).$
The congruence condition means that $\mathbf{b}(1)\equiv\mathbf{b}(2)\bmod 3$. So, the conditions on $\mathbf{b}$ are actually as follows.
$\mathbf{b}(1),\mathbf{b}(2)\in{\mathbb{Z}}_{+},\ \mathbf{b}(1)+\mathbf{b}(2)=g+2,\ \mathbf{b}(1)\equiv\mathbf{b}(2)\bmod 3.$
The list (and number) of corresponding strongly admissible functions $\mathbf{a}$ depends on $g\bmod 3$. Namely, there are the natural three cases.
  • (i)
    $g\equiv 1\bmod 3$, i.e., $g=3k+1$ where $k$ is a nonnegative integer. Then
    $\mathbf{b}(1)+\mathbf{b}(2)=g+2=3k+3=3(k+1),$
    and therefore both $\mathbf{b}(1)$ and $\mathbf{b}(2)$ are divisible by $3$. Hence there are exactly $(k+2)$ options for $\mathbf{b}$, namely,
    (35) $\mathbf{b}(1)=3d,\ \mathbf{b}(2)=3(k+1-d);\ d=0,\dots,(k+1).$
    The corresponding strongly admissible $\mathbf{a}$ are as follows (where $d=0,\dots,(k+1)$).
    $\mathbf{a}(2)=\frac{1}{3}\left(\mathbf{b}(1)+2\mathbf{b}(2)\right)-1=d+2\cdot(k+1-d)-1=d+2(k+1-d)-1=(2k+1)-d;$
    $\mathbf{a}(1)=\frac{1}{3}\left(2\mathbf{b}(1)+\mathbf{b}(2)\right)-1=2d+(k+1-d)-1=k+d.$
    So, we get
    (36) $\mathbf{a}(1)=k+d,\ \mathbf{a}(2)=(2k+1)-d;\ d=0,\dots,k+1.$
    The number of strongly admissible $\mathbf{a}$’s is
    $k+2=\frac{g+5}{3}.$
    Clearly, if $g=1$ then all both admissible functions are strongly admissible. On the other hand, if $g>1$ then there are precisely (2g2)/3(2g-2)/3 admissible functions that are not strongly admissible.
  • (ii)
    $g\equiv 2\bmod 3$, i.e., $g=3k+2$ where $k$ is a nonnegative integer. Then
    $\mathbf{b}(1)+\mathbf{b}(2)=g+2=3k+4=3(k+1)+1,$
    and therefore both $\mathbf{b}(1)-2$ and $\mathbf{b}(2)-2$ are divisible by $3$. Hence there are exactly $(k+1)$ options for $\mathbf{b}$, namely,
    (37) $\mathbf{b}(1)=3d+2,\ \mathbf{b}(2)=3(k-d)+2;\ (d=0,\dots,k).$
    The corresponding strongly admissible $\mathbf{a}$ are as follows (where $d=0,\dots,k$).
    $\mathbf{a}(2)=\frac{1}{3}\left(\mathbf{b}(1)+2\mathbf{b}(2)\right)-1=d+2(k-d)+2-1=(2k+1)-d;$
    $\mathbf{a}(1)=\frac{1}{3}\left(2\mathbf{b}(1)+\mathbf{b}(2)\right)-1=2d+(k-d)+2-1=(k+1)+d.$
    So, we get
    (38) $\mathbf{a}(1)=(k+1)+d,\ \mathbf{a}(2)=(2k+1)-d;\ d=0,\dots,k.$
    The number of strongly admissible $\mathbf{a}$’s is
    $k+1=\frac{g+1}{3}.$
    So, there are precisely $2(g+1)/3$ admissible functions that are not strongly admissible.
  • (iii)
    $g\equiv 0\bmod 3$, i.e., $g=3k$ where $k$ is a positive integer. Then
    $\mathbf{b}(1)+\mathbf{b}(2)=g+2=3k+2,$
    and therefore both $\mathbf{b}(1)-1$ and $\mathbf{b}(2)-1$ are divisible by $3$. Hence there are exactly $(k+1)$ options for $\mathbf{b}$, namely,
    (39) $\mathbf{b}(1)=3d+1,\ \mathbf{b}(2)=3(k-d)+1;\ (d=0,\dots,k).$
    The corresponding $\mathbf{a}$ are as follows (where $d=0,\dots,k$).
    $\mathbf{a}(2)=\frac{1}{3}\left(\mathbf{b}(1)+2\mathbf{b}(2)\right)-1=d+2(k-d)+1-1=2k-d;$
    $\mathbf{a}(1)=\frac{1}{3}\left(2\mathbf{b}(1)+\mathbf{b}(2)\right)-1=2d+(k-d)+1-1=k+d.$
    So, we get
    (40) $\mathbf{a}(1)=k+d,\ \mathbf{a}(2)=2k-d;\ d=0,\dots,k.$
    The number of strongly admissible $\mathbf{a}$’s is
    $k+1=\frac{g+3}{3}.$
    So, there are precisely $2g/3$ admissible functions that are not strongly admissible.
5. The CM case
We use the notation and assumptions of Section 2. Suppose that $\mathrm{dim}(X)=g=(p-1)/2$, i.e. 2g=(p1)2g=(p-1). Then $X$ becomes an abelian variety of CM type with multiplication by the CM field ${\mathbb{Q}}(\zeta_{p})$ of degree (p1)(p-1). The corresponding nonnegative multiplicity function $\mathbf{a}_{X}=\mathbf{a}_{X,\delta}$ enjoys the property
$\mathbf{a}_{X,\delta}(h)+\mathbf{a}_{X,\delta}(-h)=\frac{2g}{p-1}=1,$
which means that for each $h\in G$ either
$\mathbf{a}_{X,\delta}(h)=1,\quad\mathbf{a}_{X,\delta}(-h)=0$
or
$\mathbf{a}_{X,\delta}(h)=0,\quad\mathbf{a}_{X,\delta}(-h)=1.$
To each $h\in G=({\mathbb{Z}}/p{\mathbb{Z}})^{*}$ corresponds the field embedding
$\psi_{h}:{\mathbb{Q}}(\zeta_{p})\to{\mathbb{C}},\ \zeta_{p}\mapsto\zeta_{p}^{h}.$
Clearly, the CM type of $X$ is the (p1)/2(p-1)/2-element set
$\Psi=\Psi_{X,\delta}=\{\phi_{h}\mid\mathbf{a}_{X,\delta}(h)=1\}.$
Example 5.1.
Let $\mathcal{C}$ be the smooth projective model of the plane affine curve $y^{2}=1-x^{p}$. Then $\mathcal{C}$ has genus g=(p1)/2g=(p-1)/2 and admits an automorphism $\delta_{\mathcal{C}}:\mathcal{C}\to\mathcal{C}$ induced by
$(x,y)\mapsto(\zeta_{p}x,y).$
Let $\mathcal{J},\lambda)$ be canonically principally polarized jacobian of $\mathcal{C}$ endowed by the automorphism $\delta\in\mathrm{Aut}(\mathcal{J},\lambda)$ induced by $\delta_{\mathcal{C}}$. It is known [13, Example 15.4 (2)] that $\zeta_{p}\to\delta$ can be extended to the ring homomorphism ${\mathbb{Z}}[\zeta_{p}]\to\mathrm{End}(\mathcal{J})$ (i.e., $\sum_{i=0}^{p-1}\delta^{i}=0$), which makes $\mathcal{J}$ an abelian variety of CM type with multiplication by ${\mathbb{Q}}(\zeta_{p})$ and its CM type $\Psi$ is $\{\psi_{i}\mid 1\leq i\leq g=(p-1)/2\}$. This means that the corresponding multiplicity function $\mathbf{a}_{\mathcal{J},\delta}(h)$ is as follows.
$\mathbf{a}_{\mathcal{J},\delta}(i\bmod p)=1\ \text{ if }1\leq i\leq(p-1)/2;$
$\mathbf{a}_{\mathcal{J},\delta}(i\bmod p)=0\ \text{ if }i\geq(p-1)/2.$
Recall that $p$ is an odd prime. The case $p=3$ (with arbitrary $g$) was discussed in detail in Section 4. The following assertion deals with $p>3$ when g=(p1)/2g=(p-1)/2.
Theorem 5.2.
Let $p>3$ and g=(p1)/2g=(p-1)/2.
Then the number of strongly admissible functions of degree $g$ is $(p^{2}-1)/6$. In particular, every well rounded function of degree $g$ is strongly admissible if and only if $p\in\{5,7\}$.
Proof.
So, we need to compute the number of strongly admissible functions when 2g=p12g=p-1. By Theorem 2.4 each strongly admissible function is of the form $\frac{(p-1)}{p}\cdot\mathbf{b}*\mathbf{j}(-v)-1$ where the nonnegative integer-valued function $\mathbf{b}:G\to{\mathbb{Z}}_{+}$ enjoys the properties
(41) $\sum_{h\in G}\mathbf{b}(h)=\frac{2g}{p-1}+2=1+2=3;$
(42) $\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})\in p{\mathbb{Z}}.$
Taking into account that
$1\leq\mathbf{j}(h^{-1})<p\quad\forall h\in G,$
we conclude that
$\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})<\left(\sum_{h\in G}\mathbf{b}(h)\right)p=3p.$
It follows that (42) means that either
(43) $\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})=p$
or
(44) $\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})=2p.$
Let us compute the number of functions $\mathbf{b}$ that enjoy either properties (41) and (43) or properties (41) and (44). First, let us prove that
(45) $\mathbf{b}(h)\in\{0,1,2\}\ \forall h\in G.$
Indeed, it follows readily from (41) that
$\mathbf{b}(h)\in\{0,1,2,3\}\ \forall h\in G.$
Suppose that $\mathbf{b}(v)=3$ for some $v\in G$. Then it follows from (41) that all other values of $\mathbf{b}$ are zeros. Now (42) implies that
$3\cdot\mathbf{j}(v^{-1})=\mathbf{b}(v)\mathbf{j}(v^{-1})=\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})\in p{\mathbb{Z}}.$
Since a positive integer $\mathbf{j}(v^{-1})$ is strictly less than $p$, the prime $3$ must divide the prime $p$, which is not true, because $p\neq 3$. The obtained contradiction proves (45).
Now notice that
(46) $\mathbf{j}(v)+\mathbf{j}(-v)=p\ \forall v\in G$
(it follows readily from the very definition of $\mathbf{j}$). Let us consider the function
(47) $\bar{\mathbf{b}}:G\to{\mathbb{Z}}_{+},\ h\mapsto\mathbf{b}(-h).$
Clearly,
$\sum_{h\in G}\bar{\mathbf{b}}(h)=\sum_{h\in G}\mathbf{b}(h)=3;\quad\bar{\mathbf{b}}(G)=\mathbf{b}(G)\subset\{0,1,2\}.$
On the other hand,
$\sum_{h\in G}\bar{\mathbf{b}}(h)\mathbf{j}(h^{-1})=\sum_{h\in G}\mathbf{b}(-h)\mathbf{j}(h^{-1})=\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(-h^{-1})=\sum_{h\in G}\mathbf{b}(h)\left(p-\mathbf{j}(h^{-1})\right)=$
$p\left(\sum_{h\in G}\mathbf{b}(h)\right)-\left(\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})\right)=3p-\left(\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})\right).$
It follows that
$\sum_{h\in G}\bar{\mathbf{b}}(h)\mathbf{j}(h^{-1})=3p-p=2p$
if $\mathbf{b}$ satisfies (43). On the other hand, if $\mathbf{b}$ satisfies (44) then
$\sum_{h\in G}\bar{\mathbf{b}}(h)\mathbf{j}(h^{-1})=3p-2p=p.$
It follows from Theorem 3.1 combined with Remark 2.2(iii) that
$h\mapsto\bar{\mathbf{a}}(h)=\mathbf{a}(-h)=\frac{2g}{p-1}-\mathbf{a}(h)=1-\mathbf{a}(h)$
is a strongly admissible function of degree g=(p1)/2g=(p-1)/2. This implies that $\mathbf{a}(1\bmod p)=1$ (resp. 0) if and only if $\bar{\mathbf{a}}(1\bmod p)=0$ (resp. $1$).
Notice that
$\mathbf{a}(-1\bmod p)=\frac{\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})}{p}-1.$
This implies that $\mathbf{a}(1\bmod p)=1$ (resp. 0) if and only if $\sum_{h\in G}\mathbf{b}(h)\mathbf{j}(h^{-1})=p$ (resp. $2p$).
We will need the following two auxiliary assertions.
Lemma 5.3.
Let $Q_{3}(p)$ be the set of partitions in 3 parts of $p$.
There is a natural bijection between $Q_{3}(p)$ and the set of strongly admissible functions $\mathbf{a}:G\to{\mathbb{Z}}_{+}$ of degree g=(p1)/2g=(p-1)/2 such that
$\mathbf{a}(1\bmod p)=1.$
Lemma 5.4.
The cardinality of $Q_{3}(p)$ is $(p^{2}-1)/12$.
End of proof of Theorem 5.2 (modulo Lemma 5.3). The map $\mathbf{a}\mapsto\bar{\mathbf{a}}$ is a bijection between the sets of strongly admissible functions that take on at $1\bmod p$ the the values 0 and $1$ respectively. This implies that the number of of strongly admissible functions functions of degree (p1)/2(p-1)/2 is twice the cardinality of $Q_{3}(p)$, i.e., this number is
$2\cdot\frac{p^{2}-1}{12}=\frac{p^{2}-1}{6}.$
which proves first assertion of Theorem 5.2. In order to prove the second one, recall the the number of well rounded functions of degree (p1)/2(p-1)/2 is $2^{(}p-1)/2$. It remains to notice that $2^{(p-1)/2}>(p^{2}-1)/6$ if $p\geq 11$ while $2^{(p-1)/2}=(p^{2}-1)/6$ if $p\in\{5,7\}$. ∎
Proof of Lemma 5.3.
If $v\in G$ then we write $\mathbf{\delta}_{v}:G\to{\mathbb{Z}}_{+}$ for the corresponding delta function that takes on value $1$ at $v$ and vanishes elsewhere.
Each element $M$ of $Q_{3}(p)$ may be viewed as a non-ordered triple $\{m_{1},m_{2},m_{3}\}$ of positive integers, whose sum $\sum_{i=1}^{3}m_{i}=p$. Clearly,
$p>m_{i}\neq p-m_{j}\quad\forall i,j\in\{1,2,3\}.$
Let us define
$h_{i}:=(m_{i}\bmod p)^{-1}\in({\mathbb{Z}}/p{\mathbb{Z}})^{*}=G.$
Clearly,
(48) $h_{i}^{-1}=m_{i}\bmod p,\ \ \mathbf{j}(h_{i}^{-1})=m_{i},\ \ h_{i}^{-1}\neq h_{j}^{-1}\quad\forall i,j=1,2,3.$
Let us consider the function
$\mathbf{b}_{M}=\sum_{i=1}^{3}\mathbf{\delta}_{h_{i}}:G\to{\mathbb{Z}}_{+}.$
Clearly,
$\sum_{h\in G}\mathbf{b}_{M}(h)=\sum_{i=1}^{3}\left(\sum_{h\in G}\mathbf{\delta}_{h_{i}}(h)\right)=\sum_{i=1}^{3}1=3,$
$\sum_{h\in G}\mathbf{b}_{M}(h)\mathbf{j}(h^{-1})=\sum_{i=1}^{3}\left(\sum_{h\in G}\mathbf{\delta}_{h_{i}}(h)\mathbf{j}(h^{-1})\right)=\sum_{i=1}^{3}\mathbf{j}(h_{i}^{-1})=\sum_{i=1}^{3}m_{i}=p.$
By Theorem 3.1, the function
$\mathbf{a}_{M}:G\to{\mathbb{Z}}_{+},\ \ v\mapsto\frac{p-1}{p}b_{M}*\mathbf{j}(-v)-1=\frac{\sum_{h\in G}b_{M}(h)\mathbf{j}(-h^{-1}v)}{p}-1$
is strongly admissible; in addition,
$\mathbf{a}_{M}(1\bmod p)=\frac{2g}{p-1}-\mathbf{a}_{M}(-1\bmod p)=1-\left(\frac{\sum_{h\in G}b_{M}(h)\mathbf{j}(h^{-1})}{p}-1\right)=1-\left(\frac{p}{p}-1\right)=1.$
Let us consider the map
$\mathcal{T}:M\to\mathbf{a}_{M}$
from $Q_{3}(p)$ to the set of strongly admissible functions $G\to{\mathbb{Z}}_{+}$ of degree (p1)/2(p-1)/2 that take on value $1$ at $1\bmod p$. Let us check that $\mathcal{T}$ is bijective.
In order to check the injectiveness of $\mathcal{T}$, notice that in light of the last inequality of (48)
$\mathbf{b}_{M}(h)=\max\{\mathbf{b}_{M}(h)-\mathbf{b}_{M}(-h),\ 0\},$
which means that the function $\mathbf{b}_{M}(h)$ is uniquely determined by its “odd part”. It follows from Corollary 2.14 that $\mathcal{T}$ is injective.
In order to check that $\mathcal{T}$ is surjective, let us start with a strongly admissible function $\mathbf{a}:G\to{\mathbb{Z}}_{+}$ of degree (p1)/2(p-1)/2 with $\mathbf{a}(1\bmod p)=1$. We know that there is a function
$\mathbf{b}:G\to\{0,1,2\}\subset{\mathbb{Z}}_{+}$
such that
(49) $\sum_{h\in G}\mathbf{b}(h)=3,\quad\sum_{h\in G}b(h)\mathbf{j}(h^{-1})=p;$
$\mathbf{a}(v)=\frac{p-1}{p}\mathbf{b}*\mathbf{j}(-v)-1=\frac{\sum_{h\in G}b(h)\mathbf{j}(-h^{-1}v)}{p}-1\ \forall v\in G.$
We need to find a partition $M$ of $p$ in three parts such that $\mathbf{b}=\mathbf{b}_{M}$ (which would imply that that $\mathbf{a}=\mathbf{a}_{M}$). In light of of first equality of (49), there are precisely two following cases.
  • Case 1
    There is $h_{2}\in G$ such that $\mathbf{b}(h_{2})=2$. Such an $h_{2}$ is obviously unique. There is precisely one element of $G$, say, $h_{1}$, such that $\mathbf{b}(h_{1})=1$. The function $\mathbf{b}$ vanishes at every other element of $G$. This means that
    $\mathbf{b}=2\mathbf{\delta}_{h_{2}}+\mathbf{\delta}_{h_{1}}=\mathbf{\delta}_{h_{2}}+\mathbf{\delta}_{h_{2}}+\mathbf{\delta}_{h_{1}}.$
  • Case 2
    There are precisely three elements of $G$, say, $h_{1},h_{2},h_{3}$ such that
    $\mathbf{b}(h_{i})=1\quad\forall i=1,2,3.$
    The function $\mathbf{b}$ vanishes at every other element of $G$. This means that
    $\mathbf{b}=\mathbf{\delta}_{h_{1}}+\mathbf{\delta}_{h_{2}}+\mathbf{\delta}_{h_{3}}.$
In Case 1 let us put
$M:=\{m_{1}=\mathbf{j}(h_{2}^{-1}),\ m_{2}=\mathbf{j}(h_{2}^{-1}),\ m_{3}=\mathbf{j}(h_{1}^{-1})\}.$
In light of second equality of (49),
$p=\sum_{h\in G}b(h)\mathbf{j}(h^{-1})=2\mathbf{j}(h_{2}^{-1})+\mathbf{j}(h_{1}^{-1})=m_{1}+m_{2}+m_{3},$
i.e., $M$ is a partition of $M$ in three parts.
In Case 2 let us put
$M:=\{m_{1}=\mathbf{j}(h_{1}^{-1}),\ m_{2}=\mathbf{j}(h_{2}^{-1}),\ m_{3}=\mathbf{j}(h_{3}^{-1})\}.$
n light of second equality of (49),
$p=\sum_{h\in G}b(h)\mathbf{j}(h^{-1})=\sum_{i=1}^{3}\mathbf{j}(h_{i}^{-1})=m_{1}+m_{2}+m_{3},$
i.e., $M$ is a partition of $M$ in three parts.
In both cases we have
$\mathbf{b}=\mathbf{b}_{M}$
and therefore $\mathbf{a}=\mathbf{a}_{M}$, which proves the surjectiveness of $\mathcal{T}$.
Proof of Lemma 5.4.
We know that the prime $p>3$ is congruent to $\pm 1$ modulo $6$. This implies a well known assertion that that $p^{2}-1$ is divisible by $12$ (and even by $24$). This implies that $(p^{2}-1)/12$ is the nearest integer to $p^{2}/12$.
It is well known [1, Sect. 3.1, p. 16] that the cardinality $(Q_{3}(p))$ of $Q_{3}(p)$ coincides with the number of partitions of p3p-3 in at most three parts. On the other hand, it is known [1, Sect. 6.2, p. 58] that the latter number is the nearest integer to $\frac{((p-3)+3)^{2}}{12}$, i.e., is the nearest integer to $p^{2}/12$, which (as we have already seen) is $(p^{2}-1)/12$. This ends the proof of Lemma 5.4. ∎
6. Examples and Complements
6.1.
Let us start by recalling some generalitiies about Rosati (anti)involutions.
Let $X$ be a positive-dimensional abelian variety over an algebraically closed field, $\mathrm{End}(X)$ its endomorphism ring and ${\mathcal{Z}}_{X}$ the center of of $\mathrm{End}(X)$. Then ${\mathcal{Z}}_{X}^{0}={\mathcal{Z}}_{X}\otimes{\mathbb{Q}}$ is the center of the semisimple endomorphism algebra $\mathrm{End}^{0}(X)=\mathrm{End}(X)\otimes{\mathbb{Q}}$; the latter is a finite-dimensional semisimple $\mathbb{Q}$}-algebra and its centa ${\mathcal{Z}}_{X}$ is a direct sum of number fields, each of which is a totally real or a CM field. One may view $\mathrm{End}(X)$ and ${\mathcal{Z}}_{X}$ as orders in $\mathrm{End}^{0}(X)$ and ${\mathcal{Z}}_{X}^{0}$ respectively. If $\lambda$ is a polarization on $X$ then it gives rise to a so called Rosati involution
$\mathrm{End}^{0}(X)\to\mathrm{End}(X),\ u\mapsto u^{*},$
that sends ${\mathcal{Z}}_{X}$ to ${\mathcal{Z}}_{X}$, defining an automorphism (involution)
(50) ${\mathcal{Z}}_{X}^{0}\to{\mathcal{Z}}_{X}^{0},\ u\to u^{*}$
of the semisimple commutative $\mathbb{Q}$}-algebra ${\mathcal{Z}}_{X}$. It is known that this involution respects the splitting of ${\mathcal{Z}}_{X}^{0}$ into the direct sum of fields. In addition, it acts as the identity map on each totally real factor of ${\mathcal{Z}}_{X}$ and as the complex conjugation on each CM-factor. This implies the $\mathbb{Q}$}-algebra automorphism (50) does not depend on the choice of $\lambda$.
On the other hand, if $u$ is automorphism of $X$ then $u\in\mathrm{Aut}(X,\lambda)$ if and only if $u^{*}u=1_{X}$. This implies that if $u\in{\mathcal{Z}}_{X}$ respects one polarization on $X$ then it respects all of them!
Theorem 6.2.
Let $p$ be an odd prime, $g$ a positive unteger such that (p1)(p-1) divides $2g$, $X$ a complex $g$-dimensional abelian variety endowed with the ring embedding
$\kappa{\mathbb{Z}}[\zeta_{p}]:\hookrightarrow{\mathcal{Z}}_{X}\subset\mathrm{End}(X),\ 1\mapsto 1_{X}.$
Let us put
$\delta:=\kappa(\zeta_{p})\in{\mathcal{Z}}_{X}\subset\mathrm{End}(X).$
If $X$ is isomorphic as an algebraic variety to the jacobian of a smooth connected projective curve of genus $g$ then it enjoys one of the following two properties.
  • (i)
    $\frac{2g}{p-1}\leq p-2.$
  • (ii)
    Every primitive $p$th root of unity $\zeta$ is an eigenvalue of $\delta_{\Omega}:\Omega^{1}(X)\to\Omega^{1}(X)$ and its multiplicity is greater or equal than
    $\frac{1}{p}\cdot\frac{2g}{p-1}-\frac{p-2}{p}.$
Proof.
Clearly, $\kappa$ extends by $\mathbb{Q}$}-linearity to the embedding of $\mathbb{Q}$}-algebras
${\mathbb{Q}}(\zeta_{p}):\hookrightarrow{\mathcal{Z}}_{X}\subset\mathrm{End}^{0}(X),\quad 1\mapsto 1_{X},\ \zeta_{p}\mapsto\delta,$
which we continue to denote by $\kappa$. Since ${\mathbb{Q}}(\zeta_{p})$ is a CM field, the center ${\mathcal{Z}}_{X}$ is either a CM field or a product of CM fields, and the Rosati involution coincides with the complex conjugation on each factor. It follows that
$\kappa(\bar{u})=\left(\kappa(u)\right)^{*}\quad\forall u\in{\mathbb{Q}}(\zeta_{p}).$
Taking into account that $\bar{\zeta_{p}}\zeta_{p}=1$, we conclude that
$\delta^{*}\delta=\kappa\left(\bar{\zeta_{p}}\right)\kappa(\zeta_{p})=\kappa\left(\bar{\zeta_{p}}\zeta_{p}\right)=\kappa(1)=1_{X},$
i.e., $\delta^{*}\delta=1_{X}$, which means that $\delta\in\mathrm{Aut}(X,\lambda)$ for any polarization $\lambda$ on $X$. Now the desired result follows from Theorem 2.17. ∎
Corollary 6.3.
Let $p$ be an odd prime, $g_{0}$ a positive unteger such that (p1)(p-1) divides $2g_{0}$. Let $Y$ be a complex $g_{0}$-dimensional abelian variety endowed with the ring embedding
$\kappa:{\mathbb{Z}}[\zeta_{p}]\hookrightarrow{\mathcal{Z}}_{Y}\subset\mathrm{End}(Y),\ 1\mapsto 1_{Y}.$
Let us put
$\delta_{Y}:=\kappa(\zeta_{p})\in{\mathcal{Z}}_{Y}\subset\mathrm{End}(Y).$
Suppose that there is a primitive $p$th root of unity $\zeta$ that enjoys one of the following two properties.
  • (i)
    $\zeta$ is not an eigenvalue of $\delta_{Y,\Omega}:\Omega^{1}(X)\to\Omega^{1}(X)$.
  • $\zeta$ is an eigenvalue of $\delta_{Y,\Omega}$ but its multiplicity $a$ is strictly less than
    $\frac{1}{p}\cdot\frac{2g_{0}}{p-1}.$
Then the self-product $Y^{r}$ of $Y$ is not isomorphic as an algebraic variety to the jacobian of a smooth connected projective curve for all positive integers
$r>M:=\frac{p-2}{\frac{2g_{0}}{p-1}-a}.$
(In the case (i) we put $a=0$.)
Proof.
Let $r$ be a positive integer and $X=Y^{r}$. Then $g:=\mathrm{dim}(Y)=rg_{0}$, and the endomorphism ring $\mathrm{End}(X)=\mathrm{End}(Y^{r})$ is canonically isomorphic to the matrix ring $\mathrm{Mat}_{r}(\mathrm{End}(Y)$ of size $r$ over $\mathrm{End}(Y)$ with the same center as $\mathrm{End}(Y)$. In particular, the image of the diagonal embedding
$\kappa_{k}:{\mathbb{Z}}[\zeta_{p}]\hookrightarrow\oplus_{i=1}^{k}\mathrm{End}(Y)\subset\mathrm{Mat}_{r}(\mathrm{End}(Y))$
lies in the center ${\mathcal{Z}}_{Y}$. On the other hand,
$\Omega^{1}(X)=\Omega^{1}(Y^{r})=\oplus_{i=1}^{r}\Omega^{1}(Y)$
and the linear operator enjoys the following properties.
Its spectrum coincides with the spectrum of $\delta_{Y,\Omega}$. In addition, if an eigenvalue $\gamma$ of $\delta_{Y,\Omega}$ has multiplicity $a$ then it has multiplicity $ka$, viewed as an eigenvalue of $\delta_{X,\Omega}$.
Assume now that $X$ is isomorphic to the jacobian. If $\zeta$ is not an eigenvalue of $\delta_{X,\Omega}$ then it is not an eigenvalue of $\delta_{Y,\Omega}$ and therefore
$\frac{2rg_{0}}{p-1}=\frac{2g}{p-1}\leq p-2.$
This implies that
$r\leq\frac{p-2}{\frac{2rg_{0}}{p-1}}=M,$
because in this case $a=0$.
Now suppose that
7. Examples and Complements
7.1.
Let us start by recalling some generalitiies about Rosati (anti)involutions.
Let $X$ be a positive-dimensional abelian variety over an algebraically closed field, $\mathrm{End}(X)$ its endomorphism ring and ${\mathcal{Z}}_{X}$ the center of of $\mathrm{End}(X)$. Then ${\mathcal{Z}}_{X}^{0}={\mathcal{Z}}_{X}\otimes{\mathbb{Q}}$ is the center of the semisimple endomorphism algebra $\mathrm{End}^{0}(X)=\mathrm{End}(X)\otimes{\mathbb{Q}}$; the latter is a finite-dimensional semisimple $\mathbb{Q}$}-algebra and its centa ${\mathcal{Z}}_{X}$ is a direct sum of number fields, each of which is a totally real or a CM field. One may view $\mathrm{End}(X)$ and ${\mathcal{Z}}_{X}$ as orders in $\mathrm{End}^{0}(X)$ and ${\mathcal{Z}}_{X}^{0}$ respectively. If $\lambda$ is a polarization on $X$ then it gives rise to a so called Rosati involution
$\mathrm{End}^{0}(X)\to\mathrm{End}(X),\ u\mapsto u^{*},$
that sends ${\mathcal{Z}}_{X}$ to ${\mathcal{Z}}_{X}$, defining an automorphism (involution)
(51) ${\mathcal{Z}}_{X}^{0}\to{\mathcal{Z}}_{X}^{0},\ u\to u^{*}$
of the semisimple commutative $\mathbb{Q}$}-algebra ${\mathcal{Z}}_{X}$. It is known that this involution respects the splitting of ${\mathcal{Z}}_{X}^{0}$ into the direct sum of fields. In addition, it acts as the identity map on each totally real factor of ${\mathcal{Z}}_{X}$ and as the complex conjugation on each CM-factor. This implies the $\mathbb{Q}$}-algebra automorphism (51) does not depend on the choice of $\lambda$.
On the other hand, if $u$ is automorphism of $X$ then $u\in\mathrm{Aut}(X,\lambda)$ if and only if $u^{*}u=1_{X}$. This implies that if $u\in{\mathcal{Z}}_{X}$ respects one polarization on $X$ then it respects all of them!
Theorem 7.2.
Let $p$ be an odd prime, $g$ a positive unteger such that (p1)(p-1) divides $2g$, $X$ a complex $g$-dimensional abelian variety endowed with the ring embedding
$\kappa:{\mathbb{Z}}[\zeta_{p}]:\hookrightarrow{\mathcal{Z}}_{X}\subset\mathrm{End}(X),\ 1\mapsto 1_{X}.$
Let us put
$\delta:=\kappa(\zeta_{p})\in{\mathcal{Z}}_{X}\subset\mathrm{End}(X).$
If $X$ is isomorphic as an algebraic variety to the jacobian of a smooth connected projective curve of genus $g$ then it enjoys one of the following two properties.
  • (i)
    $\frac{2g}{p-1}\leq p-2.$
  • (ii)
    Every primitive $p$th root of unity $\zeta$ is an eigenvalue of $\delta_{\Omega}:\Omega^{1}(X)\to\Omega^{1}(X)$ and its multiplicity is greater or equal than
    $\frac{1}{p}\cdot\frac{2g}{p-1}-\frac{p-2}{p}.$
Proof.
Clearly, $\kappa$ extends by $\mathbb{Q}$}-linearity to the embedding of $\mathbb{Q}$}-algebras
${\mathbb{Q}}(\zeta_{p}):\hookrightarrow{\mathcal{Z}}_{X}\subset\mathrm{End}^{0}(X),\quad 1\mapsto 1_{X},\ \zeta_{p}\mapsto\delta,$
which we continue to denote by $\kappa$. Since ${\mathbb{Q}}(\zeta_{p})$ is a CM field, the center ${\mathcal{Z}}_{X}$ is either a CM field or a product of CM fields, and the Rosati involution coincides with the complex conjugation on each factor. It follows that
$\kappa(\bar{u})=\left(\kappa(u)\right)^{*}\quad\forall u\in{\mathbb{Q}}(\zeta_{p}).$
Taking into account that $\bar{\zeta_{p}}\zeta_{p}=1$, we conclude that
$\delta^{*}\delta=\kappa\left(\bar{\zeta_{p}}\right)\kappa(\zeta_{p})=\kappa\left(\bar{\zeta_{p}}\zeta_{p}\right)=\kappa(1)=1_{X},$
i.e., $\delta^{*}\delta=1_{X}$, which means that $\delta\in\mathrm{Aut}(X,\lambda)$ for any polarization $\lambda$ on $X$. Now the desired result follows from Theorem 2.17. ∎
Corollary 7.3.
Let $p$ be an odd prime, $g_{0}$ a positive unteger such that (p1)(p-1) divides $2g_{0}$. Let $Y$ be a complex $g_{0}$-dimensional abelian variety endowed with the ring embedding
$\kappa:{\mathbb{Z}}[\zeta_{p}]\hookrightarrow{\mathcal{Z}}_{Y}\subset\mathrm{End}(Y),\ 1\mapsto 1_{Y}.$
Let us put
$\delta_{Y}:=\kappa(\zeta_{p})\in{\mathcal{Z}}_{Y}\subset\mathrm{End}(Y).$
Suppose that there is a primitive $p$th root of unity $\zeta$ that enjoys one of the following two properties.
  • (i)
    $\zeta$ is not an eigenvalue of $\delta_{Y,\Omega}:\Omega^{1}(X)\to\Omega^{1}(X)$.
  • $\zeta$ is an eigenvalue of $\delta_{Y,\Omega}$ but its multiplicity $a$ is strictly less than
    $\frac{1}{p}\cdot\frac{2g_{0}}{p-1}.$
Then the self-product $Y^{r}$ of $Y$ is not isomorphic as an algebraic variety to the jacobian of a smooth connected projective curve for all positive integers
$r>M:=\frac{p-2}{\frac{1}{p}\frac{2g_{0}}{p-1}-a}.$
(In the case (i) we put $a=0$.)
Proof.
Let $r>M$ be a positive integer and $X=Y^{r}$. Then $g:=\mathrm{dim}(Y)=rg_{0}$, and the endomorphism ring $\mathrm{End}(X)=\mathrm{End}(Y^{r})$ is canonically isomorphic to the matrix ring $\mathrm{Mat}_{r}(\mathrm{End}(Y)$ of size $r$ over $\mathrm{End}(Y)$ with the same center as $\mathrm{End}(Y)$. In particular, the image of the diagonal embedding
$\kappa_{k}:{\mathbb{Z}}[\zeta_{p}]\hookrightarrow\oplus_{i=1}^{k}\mathrm{End}(Y)\subset\mathrm{Mat}_{r}(\mathrm{End}(Y)),\quad u\mapsto(\kappa(u),\dots,\kappa(u))\ \ (k\text{ times)}$
lies in the center ${\mathcal{Z}}_{Y}$. On the other hand,
$\Omega^{1}(X)=\Omega^{1}(Y^{r})=\oplus_{i=1}^{r}\Omega^{1}(Y)$
and the linear operator enjoys the following properties.
Its spectrum coincides with the spectrum of $\delta_{Y,\Omega}$. In addition, if an eigenvalue $\gamma$ of $\delta_{Y,\Omega}$ has multiplicity $a$ then it has multiplicity $ka$, viewed as an eigenvalue of $\delta_{X,\Omega}$.
Assume now that $X$ is isomorphic to the jacobian of a smooth connected projective curve. If $\zeta$ is not an eigenvalue of $\delta_{X,\Omega}$ then it is not an eigenvalue of $\delta_{Y,\Omega}$ and therefore
$\frac{2rg_{0}}{p-1}=\frac{2g}{p-1}\leq p-2.$
This implies that
$r\leq\frac{p-2}{\frac{2g_{0}}{p-1}}=M,$
because in this case $a=0$. This contradicts our assumption on $r$. So, $\zeta$ is an eigenvalue of $\delta_{X,\Omega}$. Then its multiplicity equals $ra$. Taking into account that
$(\frac{1}{p}\frac{2g_{0}}{p-1}-a>0,\ \ r>M,$
we get
$ra=r\cdot\frac{1}{p}\frac{2g_{0}}{p-1}-r\cdot\left(\frac{1}{p}\frac{2g_{0}}{p-1}-a\right)<\frac{1}{p}\frac{2g_{0}r}{p-1}-M\left(\frac{1}{p}\frac{2g_{0}}{p-1}-a\right)=\frac{1}{p}\frac{2g}{p-1}-\frac{p-2}{p}.$
In other words, $\zeta$ is an eigenvalue of $\delta_{X,\Omega}$ with multiplicity that is strictly less than
$\frac{1}{p}\frac{2g}{p-1}-\frac{p-2}{p},$
which contradicts to Theorem 2.17. The obtained contradiction implies that $X$ is not isomorphic to a jacobian. ∎
Example 7.4.
Let $p$ be an odd prime and $n\geq 4$ an integer such that $p$ does not divide $n$. We have
(52) $n=ap+c;\quad a,c\in{\mathbb{Z}}_{+};1\leq c\leq p-1;\quad 0\leq c-1\leq p-2.$
Let $f(x)\in{\mathbb{C}}[x]$ be a degree $n$ polynomial without repeated roots. Let ${\mathcal{C}}_{f,p}$ be the smooth projective model of the smooth plane affine curve $y^{p}=f(x)$. The genus $g_{0}$ of ${\mathcal{C}}_{f,p}$ is (n1)(p1)/2(n-1)(p-1)/2; hence
$\frac{2g_{0}}{p-1}=n-1=ap+(c-1).$
There is an automorphism $\tilde{\delta}\in\mathrm{Aut}({\mathcal{C}}_{f,p})$ of ${\mathcal{C}}_{f,p}$ defined by
$(x,y)\mapsto(x,\zeta_{p}y).$
Let $(J({\mathcal{C}}_{f,p}),\Theta)$ be the canonically principally polarized jacobian of $C_{f,p}$. By Albanese functoriality $\tilde{\delta}$ induces the automorphism $\delta$ of $({\mathcal{J}},\Theta)$, which satisfies the $p$th cyclotomic equation. The corresponding multiplicity function is
$\mathbf{a}_{{\mathcal{J}}({\mathfrak{C}}_{f,p}),\delta}:(-k\bmod p)\mapsto[nk/p]\ \ (1\leq k\leq p-1).$
In particular, in light of (52),
(53) $\mathbf{a}_{{\mathcal{J}}({\mathcal{C}}_{f,p}),\delta}(-1\bmod p)=[n/p]=[(ap+c)/p]=a=$
$\frac{(n-1)-(c-1)}{p}=\frac{1}{p}\frac{2g_{0}}{p-1}-\frac{(c-1)}{p}\geq\frac{2g_{0}}{p-1}-\frac{p-2}{p}.$
Let us assume that
$c>1,$
i.e., $p$ does not divide n1n-1. Then either $\zeta_{p}^{-1}$ is not an eigenvalue of
$\delta_{\Omega}:\Omega^{1}({\mathcal{J}}({\mathcal{C}}_{f,p}))\to\Omega^{1}({\mathcal{J}}({\mathcal{C}}_{f,p}))$
or it is an eigenvalue but its multiplicity $a$ is strictly less than $\frac{1}{p}\frac{2g_{0}}{p-1}$. Now it follows from Corollary 7.3 that if $\delta$ is a central element of $\mathrm{End}({\mathcal{J}}({\mathcal{C}}_{f,p}))$ then ${\mathcal{J}}({\mathcal{C}}_{f,p})^{r}$ is not isomorphic as an algebraic variety to the jacobian of a smooth connected projective curve for all positive integers
$r>\frac{p-2}{\frac{1}{p}\frac{2g_{0}}{p-1}-a}=\frac{p-2}{\frac{1}{p}(n-1)-a}=\frac{p-2}{\frac{1}{p}(pa+c-1)-a}=\frac{p-2}{(c-1)/p}=\frac{p(p-2)}{c-1}.$
In other words, ${\mathcal{J}}({\mathcal{C}}_{f,p})^{r}$ is not isomorphic to a jacobian (even if one ignores polarizations) if
$r>\frac{p(p-2)}{(c-1)}.$
Theorem 7.5.
Let $p$ be an odd prime and $n\geq 4$ an integer. Suppose that $p$ does not divide n(n1)n(n-1), i.e., $n=ap+c$ where $a,c\in{\mathbb{Z}}_{+}$. Assume additionally that
$2\leq c\leq p-1.$
Let $K$ be a subfield of $K$ that contains $\zeta_{p}$. Let
$f(x)\in K[x]\subset{\mathbb{C}}[x]$
be a degree $n$ irreducible polynomial, whose Galois group $\mathrm{Gal}(f/K)$ over $K$ enjoys one of the following two properties.
  • $n\geq 4$ and $\mathrm{Gal}(f)$ coincides with the full symmetric group $\mathbf{S}_{n}$;
  • $n\geq 5$ and $\mathrm{Gal}(f)$ coincides with the alternating group $\mathbf{A}_{n}$.
Then ${\mathcal{J}}({\mathcal{C}}_{f,p})^{r}$ is not isomorphic as an algebraic variety to the jacobian of a smooth connected projective curve for all positive integers
$r>\frac{p(p-2)}{c-1}.$
In particular, ${\mathcal{J}}({\mathcal{C}}_{f,p})^{r}$ is not isomorphic to a jacobian if
r>p(p2).r>p(p-2).
Proof.
It is known [17, 18, 21, 22] that our assumptions on $n$ and $f(x)$ imply that the endomorphism ring
$\mathrm{End}(J({\mathcal{C}}_{f,p})={\mathbb{Z}}[\delta]\cong{\mathbb{Z}}[\zeta_{p}].$
In particular, the ring $\mathrm{End}(J({\mathcal{C}}_{f,p})$ is commutative and therefore $\delta$ lies in its center. Now the desired results follow readily from considerations of Example 7.4 if we take into account that $c-1\geq 2-1=1$ and therefore
$p(p-2)\geq\frac{p(p-2)}{c-1}.$
8. CM case revisited
Theorem 8.1.
Let $p$ be an odd prime, $X$ an abelian variety of dimension (p1)/2(p-1)/2 endowed with a ring endomorphism
$\kappa:{\mathbb{Z}}[\zeta_{p}]\cong\mathrm{End}(X).$
If r>(p2)r>(p-2) is an integer then $X^{r}$ is not isomorphic as an algebraic variety to the jacobian of a smooth connected projective curve.
Proof.
Let us consider
$\delta:=\kappa(\zeta_{p})\in\mathrm{Aut}(X).$
Clearly, $\delta$ satisfies the $p$th cyclotomic equation in $\mathrm{End}(X)$. On the other hand, the complex vector space $\Omega^{1}(X)$ has dimension (p1)/2(p-1)/2 that is strictly less that (p1)(p-1). Hence, there is a primitive $p$th root of unity, say, $\zeta$ that is not an eigenvalue of $\delta_{X}:\Omega^{1}(X)\to\Omega^{1}(X)$. Now the desired result follows from Corollary 7.3 applied to
$Y=X,\ g_{0}=\frac{p-1}{2},\ a=0.$
Example 8.2.
Let $p=3$ and $X$ an elliptic curve with $\mathrm{End}(X)={\mathbb{Z}}[\zeta_{3}]$. Then $X^{r}$ is not isomorphic as an algebraic variety to the jacobian of a smooth connected projective curve. (This assertion is well known for $r=2$, see [5].)